Assume​ Evco, Inc., has a current stock price of $ 59 and will pay a $ 1.75 dividend in one​ year; its equity cost of capital is 17 % . What price must you expect Evco stock to sell for immediately after the firm pays the dividend in one year to justify its current​ price?

The expected price is..$?
Answer #1

equity cost of capital=(D1/Current price)+Growth rate

0.17=(1.75/59)+Growth rate

Growth rate=0.17-(1.75/59)

=0.140338983

Hence expected price=59*(1+Growth rate)

59*(1+0.140338983)

=$67.28